# MEEG and AT-DGNN: Advancing EEG Emotion Recognition with Music and Graph Learning

### v1

@@ III-B3 Multi-head Attention Module
Each subsequence $X_{w}$ derived from sliding window segmentation is first standardized using a layer normalization (LayerNorm) layer. For the normalized input $X_{w}^{\text{norm}}$, the MHA mechanism computes the $q_{ht}$, $k_{ht}$, and $v_{ht}$ as linear transformations:
$\begin{cases}q_{ht}=W^{Q}\text{LN}(X_{w,t}),\\ k_{ht}=W^{K}\text{LN}(X_{w,t}),\\ v_{ht}=W^{V}\text{LN}(X_{w,t}),\end{cases}$ (8)
where LN denotes layer normalization. The matrices $W^{Q}$, $W^{K}$, and $W^{V}$ belong to $\mathbb{R}^{d\times d_{H}}$, where $d_{H}$ represents the dimension of each attention head. The attention context vector $c_{ht}$ for each head is computed as a weighted sum of the values, using weights $\alpha_{htt^{\prime}}$ derived from the scaled dot-product attention mechanism.
The context vectors from all heads are concatenated and linearly transformed to yield the final output of the MHA layer:
$X_{w}^{\prime}=W^{O}[C_{1},\ldots,C_{H}]+X_{w},$ (9)
where $W^{O}\in\mathbb{R}^{d_{H}\times d}$ is the output projection matrix. This procedure effectively projects the combined outputs back to the original input dimension. Subsequently, $X_{w}^{\prime}$ is integrated with the original subsequence $X_{w}$ through a residual connection and normalized once more. This advanced MHA framework significantly enhances the model’s capacity to discern intricate temporal patterns and dependencies within EEG signals, thereby augmenting the precision of decoding activities.
@@ III-B4 Temporal Convolution
Following the MHA mechanism, the output, denoted as attn_output, is normalized with a LayerNorm layer and merged with the sliding window data through a residual connection to generate the EEG data $X_{i}\in\mathbb{R}^{c\times l}$, where $c$ is the number of EEG channels, and $l$ represents the sample length in the temporal dimension.
We employ multiple temporal convolutional kernels operating in parallel to extract dynamic temporal features. Each kernel’s output, $Z^{k}_{\text{tcn}}\in\mathbb{R}^{t\times c\times f_{k}}$, where $t$ denotes the total number of kernels and $f_{k}$ the feature length, acts as a digital filter capturing signals across various frequency bands. The addition of batch normalization ensures training stability, and ReLU activation introduces non-linearity.
The outputs from different kernels are concatenated along the feature dimension, forming a multi-scale output $Z_{i}^{\text{MS}}\in\mathbb{R}^{t\times c\times f_{k}}$:
$Z_{i}^{\text{MS}}=[Z^{1}_{\text{tcn}},\ldots,Z^{K}_{\text{tcn}}]$ (10)
This integration of the temporal convolution with the MHA mechanism captures local temporal dependencies effectively, enriching the feature representations and enhancing their discriminative capabilities.
@@ III-B5 Feature Fusion
After computing the outputs of the temporal learner for each window, denoted as $Z^{w}_{\text{tcn}}\in\mathbb{R}^{B\times C\times T_{w}}$, where $B$ is the batch size, $C$ represents the channel count, $T_{w}$ is the window’s temporal length, and $w$ indexes the window, these outputs are assembled into a four-dimensional tensor:
$Z_{\text{stacked}}=[Z^{1}_{\text{tcn}},Z^{2}_{\text{tcn}},\ldots,Z^{W}_{\text{tcn}}]\in\mathbb{R}^{B\times C\times W\times T_{w}}$ (11)
This tensor is subsequently rearranged and flattened to conform to the input requirements of a convolution layer, resulting in dimensions $(B,32,-1)$, where $32$ refers to the fixed channel count and $-1$ represents the flattened dimensions. A fusion convolution layer, utilizing a kernel size of 3 and a stride of 1, integrates the outputs across all windows. The output, $Z_{\text{fused}}\in\mathbb{R}^{B\times 32\times L}$, where $L$ is the combined length, processes the temporal features to capture global contextual information effectively. This approach allows for a more discriminative representation of features in EEG signals, which is crucial for constructing nodes in GNNs.
III-C Graph Learning
@@ III-C1 Graph Filtering Layer
For the extracted features, we adopt the method described by Ding et al.[16], which involves defining functional areas and performing local filtering. EEG data is divided into three functional areas: general region ($G_{g}$), frontal region ($G_{f}$), and hemispheric region ($G_{h}$). Each electrode is treated as a node, and the learned dynamic temporal representations are considered as node attributes. According to the 10-20 system, electrodes are grouped based on their scalp locations, as Fig 2.
@@
In the preprocessing step, EEG channels are systematically reordered within predefined groups to ensure adjacency of channels within each local graph. This reordering facilitates the subsequent application of localized graph operations. The reordering function is mathematically defined as:
$Z_{i}^{\text{reorder}}=F_{\text{reorder}}(Z_{i}^{\text{fuse}})$ (12)
Following the reordering, a graph filtering layer is employed to aggregate the transformed representations and learn patterns indicative of local brain activity. The local adjacency matrix, denoted as $A_{\text{local}}$, is initialized as a fully connected matrix where all elements are set to 1. The local graph filtering operation then applies a trainable filter matrix $W_{\text{local}}$ and a bias vector $b_{\text{local}}$ to the reordered channel representations:
$Z_{i}^{\text{filtered}}=\text{ReLU}(W_{\text{local}}\circ Z_{i}^{\text{reorder}}-b_{\text{local}})$ (13)
The output of the filtering process $Z_{i}^{\text{local}}$ within each graph is subsequently aggregated to form localized feature representations. This aggregation is performed using the defined aggregation function $F_{\text{aggregate}}$, resulting in the following vector representation for each local graph:
$Z_{i}^{\text{local}}=F_{\text{aggregate}}(Z_{i}^{\text{filtered}})=\begin{bmatrix}h_{1}^{\text{local}}\\ \vdots\\ h_{R}^{\text{local}}\end{bmatrix}$ (14)
@@ III-C2 Stacked DGNN
1:Input: EEG data $X_{i}\in\mathbb{R}^{E\times T}$, ground truth label $y$; graph definitions $G_{g},G_{f},G_{h}$; global adjacency matrix $W_{\text{global}}$
@@
6:     get $i$th temporal kernel size by Eq. 4;
7:     get $Z^{i}_{\text{temp}}$ by Eq. 5 using $X_{i}$ as input;
@@
9:get $Z_{T}$ by Eq. 6;
10:get the window size $n$ and segment the data by Eq. 7
11:get the output $X_{w}^{\prime}$ by Eq. 9 after MHA;
12:concatenate the feature dimension by Eq. 10
13:do feature fusion by Eq. 11 to get $Z_{\text{fused}}$;
14$\triangleright$ Graph Learning
15:do graph filtering and aggregation on each node by Eq. 12 - 14, get the output $Z_{i}^{\text{local}}$.
16:for $i\leftarrow 1$ to $3$ do
17:     get DGNN output by Eq. 17, Eq. 18 and Eq. 19;
@@
19:Get pred by Eq. 20;
20:Return pred
We propose a novel stacked DGNN architecture engineered to discern complex structural relationships within graphs. Each layer of this model adaptively recalculates the adjacency matrix using input features, facilitating the effective capture of hierarchical feature variations across multiple levels.
The architecture incorporates multiple GNN layers. In each layer, features are grouped and aggregated into matrices, with each group containing $k$ channels, resulting in an aggregated feature matrix $Z_{\text{agg}}$. The adjacency matrix for each layer is dynamically computed based on the feature similarity matrix $S$, defined as:
$S=Z_{\text{agg}}\cdot Z_{\text{agg}}^{T}$ (15)
To preserve node connectivity within the graph, self-loops are introduced by adding the identity matrix $I$ to $S$, and the modified adjacency matrix $A$ is subsequently normalized as follows:
@@
During forward propagation, the initial layer performs graph convolution using the normalized adjacency matrix $\tilde{A}^{(1)}$, weight matrix $W^{(1)}$, and bias vector $b^{(1)}$. The output from this layer is given by:
$H^{(1)}=\text{ReLU}(\tilde{A}^{(1)}\cdot Z_{\text{agg}}\cdot W^{(1)}+b^{(1)})$ (17)
Subsequent layers $l=2,\ldots,L-1$ process the output of the previous layer $H^{(l-1)}$ with corresponding layer-specific parameters $\tilde{A}^{(l)}$, $W^{(l)}$, and $b^{(l)}$, to produce:
$H^{(l)}=\text{ReLU}(\tilde{A}^{(l)}\cdot H^{(l-1)}\cdot W^{(l)}+b^{(l)})$ (18)
The final output $H^{(L)}$ from the last hidden layer $H^{(L-1)}$ utilizes the parameters of the ultimate layer $\tilde{A}^{(L)}$, $W^{(L)}$, and $b^{(L)}$:
$H^{(L)}=\text{ReLU}(\tilde{A}^{(L)}\cdot H^{(L-1)}\cdot W^{(L)}+b^{(L)})$ (19)
The final network output is computed through a sequence of operations including batch normalization, dropout, and softmax activation, structured as:
$\text{Output}=\Phi_{\text{softmax}}(\mathcal{F}_{\text{dropout}}(\Gamma(\mathcal{F}_{\text{bn}}(H^{(L)}))))$ (20)
@@
This structured approach not only ensures detailed feature extraction at various levels but also enhances the model’s ability to learn and generalize from complex graph structures. Finally, the procedure of AT-DGNN can be summarized in Algorithm 1.
@@ IV Experiments
TABLE I: The ACC and F1 scores (mean+std) on the MEEG dataset (trial-wise $10$-fold cross-validation)
Method Arousal Valence Arousal Valence
ACC (%) std (%) F1 (%) std (%) ACC (%) std (%) F1 (%) std (%)
EEGNet 73.40 17.82 73.03 18.31 70.00 18.06 68.52 18.95
DeepConvNet 74.99 19.91 72.04 22.86 72.38 18.87 73.03 19.27
ShallowConvNet 78.63 14.68 78.61 15.25 77.21 17.42 76.96 17.74
TSception 79.23 11.51 81.30 9.84 78.34 15.03 81.18 12.08
EEGTCNet 76.36 17.05 73.79 19.66 69.72 16.06 54.28 22.43
TCNet-Fusion 74.88 17.62 75.09 17.80 73.63 17.56 72.95 18.09
ATCNet 82.01 14.06 79.17 17.57 73.63 17.56 72.95 18.09
DGCNN 82.12 12.05 81.83 12.45 82.72 11.00 82.40 11.44
LGGNet-Fro 80.90 13.06 80.76 13.26 82.34 12.84 81.66 13.91
LGGNet-Gen 80.86 12.88 81.47 12.24 82.37 13.20 82.47 13.22
LGGNet-Hem 81.08 13.43 80.47 14.31 82.53 12.71 81.58 14.13
AT-DGNN-Fro 83.06 11.54 83.88 10.89 84.88 10.39 84.92 10.56
AT-DGNN-Gen 82.63 12.12 83.97 10.67 85.31 9.88 85.89 9.66
AT-DGNN-Hem 82.34 11.83 82.12 12.40 84.94 10.16 85.43 9.79
@@ IV-A MEEG Dataset
In studies on emotion induction, visual stimuli such as images, videos, and text are commonly employed in experiments; however, participants’ responses to these stimuli may be influenced by their cultural backgrounds. Considering the auditory cortex exhibits emotion-specific functional connections with a wide array of limbic, paralimbic, and neocortical structures, suggesting a more extensive role in emotion processing than previously understood [25], we opted to induce emotions in participants using music in the MEEG dataset. We carefully selected 20 music clips, each lasting one minute, characterized by distinct arousal and valence levels, decidedly non-neutral. To minimize emotional carryover, a 15-second pause was introduced between clips [26]. The chosen music pieces included lesser-known works of Western classical music by composers such as Shostakovich and Tchaikovsky. The study involved 32 students from Shandong University, aged between 20 and 25 years, who had no formal music education and were unfamiliar with Western classical music. Each participant signed an informed consent form prior to the study. This experimental design aimed to minimize the influence of social and cultural backgrounds on emotional responses, ensuring that changes in emotions were solely due to musical stimuli, thereby enhancing the accuracy of the experimental results. EEG data were collected using a 32-channel BCI device NeuSen.W32 from Neuracle Tech, employing the same electrode channels as in the DEAP dataset [7]. The data were sampled at a frequency of 1000 Hz. The EEG data were annotated based on the arousal and valence of the music during various stages of the experiment.

### v2

@@ III-B3 Multi-head Attention Module
Each subsequence $X_{w}$ derived from sliding window segmentation is processed as described in (1). The attention context vector $c_{ht}$ for each head is computed as a weighted sum of the values, with weights $\alpha_{htt^{\prime}}$ determined by the scaled dot-product attention mechanism as (2), and then used in (3).
The context vectors $c_{1t},\ldots,c_{Ht}$ from all heads are concatenated and linearly transformed to yield the final output of the MHA layer:
$\textit{X}_{w}^{\prime}=W^{O}[\textit{c}_{1t},\ldots,\textit{c}_{Ht}]+\textit{X}_{w},$ (9)
where $W^{O}\in\mathbb{R}^{d_{H}\times d}$ is the output projection matrix. This procedure projects the combined outputs back to the original input dimension. Subsequently, $X_{w}^{\prime}$ is integrated with the original subsequence $X_{w}$ through a residual connection and normalized once more. This advanced MHA framework significantly enhances the model’s capacity to discern intricate temporal patterns and dependencies within EEG signals, thereby augmenting the precision of decoding activities.
@@ III-B4 Temporal Convolution
Following the MHA mechanism, the output $\textit{X}_{w}^{\prime}$ is normalized using LayerNorm. The normalized EEG data are denoted as $X_{i}\in\mathbb{R}^{c\times l}$, where $c$ represents the number of EEG channels and $l$ represents the temporal dimension’s sample length.
To extract dynamic temporal features, we employ a temporal convolutional block on each window of the EEG data. The normalized EEG data $X_{i}$ serves as the input to this block. The temporal convolutional block processes $X_{i}$ and generates an output tensor $Z_{\textit{tcn}}^{w}\in\mathbb{R}^{B\times C\times T_{w}}$, where $T_{w}$ represents the window’s temporal length and $w$ indexes the window. Batch normalization is applied to the output to ensure training stability, followed by a $\textit{ReLU}(\cdot)$ activation function to introduce non-linearity. The output $Z_{\textit{tcn}}^{w}$ is computed as follows:
$Z_{\textit{tcn}}^{w}=\textit{ReLU}(\textit{BatchNorm}(\textit{Conv1d}(X_{i})))$ (10)
@@ III-B5 Feature Fusion
The outputs of the temporal learner for each window, denoted as $Z^{w}_{\textit{tcn}}\in\mathbb{R}^{B\times C\times T_{w}}$, are assembled into a four-dimensional tensor $Z_{\textit{stacked}}$:
$Z_{\textit{stacked}}=[Z^{1}_{\textit{tcn}},Z^{2}_{\textit{tcn}},\ldots,Z^{W}_{\textit{tcn}}]\in\mathbb{R}^{B\times C\times W\times T_{w}}$ (11)
This tensor is rearranged and flattened to meet the convolution layer’s input requirements, resulting in dimensions $(B,32,-1)$, where $32$ is the fixed channel count, and $-1$ represents the flattened dimensions. A fusion convolution layer with a kernel size of 3 and a stride of 1 integrates the outputs across all windows. The final output $Z_{\text{fused}}\in\mathbb{R}^{B\times 32\times L}$, where $L$ is the combined length, captures global contextual information, allowing for a more discriminative representation of features in EEG signals, which is crucial for constructing nodes in GNNs.
III-C Graph-based Learning
@@ III-C1 Graph Filtering Layer
For the extracted features, we adopt the method described by [16], which involves defining functional areas and performing local filtering. The electrodes are divided into three functional areas: the general region ($G_{g}$), the frontal region ($G_{f}$), and the hemispheric region ($G_{h}$), as shown in Fig. 2. Each channel is treated as a node, and the learned dynamic temporal representations are considered as node attributes.
@@
In the preprocessing step, EEG channels are systematically reordered within predefined groups to ensure adjacency of channels within each local graph. This reordering enhances the effectiveness of localized graph-based operations applied in subsequent stages. Mathematically, the reordering of channel representations, denoted by $Z_{\textit{reorder}}^{i}$, is defined by the function $F_{\textit{reorder}}(\cdot)$, which operates on the fused data from EEG channels at the $i$-th observation representations $Z_{\textit{fuse}}^{i}$. This relationship is expressed as:
$Z_{\textit{reorder}}^{i}=F_{\textit{reorder}}(Z_{\textit{fuse}}^{i})$ (12)
After reordering, a graph filtering layer is employed to aggregate the transformed representations of EEG channels using a fully connected local adjacency matrix $A_{\textit{local}}$, where all elements are set to 1. The filtered output $Z_{\textit{filtered}}^{i}$ is obtained by applying a trainable filter matrix $W_{\textit{local}}$ and a bias $b_{\textit{local}}$ to the reordered data, followed by the application of the $\textit{ReLU}(\cdot)$ activation function:
$Z_{\textit{filtered}}^{i}=\textit{ReLU}(W_{\textit{local}}\circ Z_{\textit{reorder}}^{i}+b_{\textit{local}})$ (13)
The element-wise product symbol $\circ$ captures localized interactions within the brain.
The output of the filtering process, $Z_{\textit{local}}^{i}$, is aggregated to form localized feature representations using the aggregation function $F_{\textit{aggregate}}(\cdot)$. The vector representation for each local graph is:
$Z_{\textit{local}}^{i}=F_{\textit{aggregate}}(Z_{\textit{filtered}}^{i})=\begin{bmatrix}h_{\textit{local}}^{1}\\ \vdots\\ h_{\textit{local}}^{R}\end{bmatrix}$ (14)
where $h_{\textit{local}}^{1},\dots,h_{\textit{local}}^{R}$ are elements of the localized feature representation vector. These features $Z_{\textit{local}}^{i}$ are used to construct the overall feature matrix $Z_{\textit{agg}}$:
$Z_{\textit{agg}}=\begin{bmatrix}Z_{\textit{local}}^{1}\\ Z_{\textit{local}}^{2}\\ \vdots\\ Z_{\textit{local}}^{11}\end{bmatrix}$ (15)
@@ III-C2 Stacked DGNN
1:Input: EEG data $X_{i}\in\mathbb{R}^{E\times T}$, ground truth label $y$
@@
6:     get $i$th temporal kernel size by (5);
7:     get $Z^{i}_{\textit{temp}}$ by (6) using $X_{i}$ as input;
@@
9:get $Z_{T}$ by (7);
10:get the window size $n$ and segment the data by (8):
11:for $w\leftarrow 1$ to $n$ do
12:     get window $X_{w}\in\mathbb{R}^{B\times E\times W}$;
13:     normalize $X_{w}$ using LayerNorm;
14:     process $X_{w}$ with MHA to get $X_{w}^{\prime}$ by (9);
15:     normalize $X_{w}^{\prime}$ using LayerNorm;
16:     apply TCN on $X_{w}^{\prime}$ to get $Z_{\textit{tcn}}^{w}$ by (10);
17:     add $Z_{\textit{tcn}}^{w}$ to $Z_{\textit{stacked}}$;
@@
19:apply Fusion Conv to $Z_{\textit{stacked}}$ to get $Z_{\textit{fused}}$ by (11);
20$\triangleright$ Graph-based Learning
21:do graph filtering and aggregation on each node by (12) - 15 to get $Z^{i}_{\textit{local}}$;
22:for $i\leftarrow 1$ to $3$ do
23:     get DGNN output by (18);
24:end for
25:get pred by (19);
26:Return pred
We introduce a stacked DGNN architecture to discern complex structural relationships within graphs. Each layer adaptively recalculates the adjacency matrix based on input features, effectively capturing hierarchical feature variations.
The architecture incorporates multiple GNN layers. In each layer, features are grouped and aggregated into matrices, resulting in an aggregated feature matrix $Z_{\textit{agg}}$. The adjacency matrix for each layer is dynamically computed based on the feature similarity matrix $S$:
$S=Z_{\textit{agg}}\cdot Z_{\textit{agg}}^{T}$ (16)
Self-loops are introduced by adding the identity matrix $I$ to $S$, and the modified adjacency matrix $A$ is normalized as:
@@
Here, $\tilde{D}$ is the diagonal degree matrix of $A$, with each diagonal element $\tilde{D}_{ii}$ being the sum of the $i$-th row of $A$.
During forward propagation, each DGNN layer processes the inputs using a normalized adjacency matrix $\tilde{A}_{i}$, a weight matrix $W_{i}$, and a bias vector $b_{i}$ for graph convolution. The outputs for each layer are calculated as:
$\begin{cases}H_{1}=\textit{ReLU}(\tilde{A}_{1}\cdot Z_{\textit{agg}}\cdot W_{1}+b_{1})\\ H_{i}=\textit{ReLU}(\tilde{A}_{i}\cdot H_{i-1}\cdot W_{i}+b_{i}),\quad i\in[2,n-1]\\ H_{n}=\textit{ReLU}(\tilde{A}_{n}\cdot H_{n-1}\cdot W_{n}+b_{n})\end{cases}$ (18)
Here, $n$ denotes the total number of layers in the DGNN, which is set to $3$ in our study.
The final Output is computed through a sequence of operations including batch normalization $\mathcal{F}_{\textit{bn}}(\cdot)$, dropout $\mathcal{F}_{\textit{dropout}}(\cdot)$, and softmax activation $\Phi_{\textit{softmax}}(\cdot)$, structured as:
$\textit{Output}=\Phi_{\textit{softmax}}(\mathcal{F}_{\textit{dropout}}(\Gamma(\mathcal{F}_{\textit{bn}}(H^{(L)}))))$ (19)
@@
Finally, the procedure of AT-DGNN can be summarized in Algorithm 1.
@@ IV-A MEEG Dataset
In studies on emotion induction, visual stimuli such as images, videos, and text are commonly employed in experiments. However, participants’ responses to these stimuli may be influenced by their cultural backgrounds [26]. Considering that the auditory cortex exhibits emotion-specific functional connections with a wide array of limbic, paralimbic, and neocortical structures, suggesting a more extensive role in emotion processing than previously understood [27], we opted to induce emotions in participants using music in the MEEG dataset. We carefully selected 20 music clips, each lasting one minute, characterized by distinct arousal and valence levels, decidedly non-neutral. To minimize emotional carryover, a 15-second pause was introduced between clips [28]. The chosen music pieces included lesser-known works of Western classical music by composers such as Shostakovich and Tchaikovsky. The study involved 32 students from Shandong University, aged between 20 and 25 years, who had no formal music education and were unfamiliar with Western classical music. Each participant signed an informed consent form prior to the study. This experimental design aimed to minimize the influence of social and cultural backgrounds on emotional responses, ensuring that changes in emotions were solely due to musical stimuli, thereby enhancing the accuracy of the experimental results. EEG data were collected using a 32-channel BCI device NeuSen.W32 from Neuracle Tech, employing the same electrode channels as in the DEAP dataset [7]. The data were sampled at a frequency of 1000 Hz. The EEG data were annotated based on the arousal and valence of the music during various stages of the experiment.
@@ IV-C Implement Details
TABLE I: Comparison of ACC and F1 scores (mean $\pm$ std) for the MEEG dataset under trial-wise $10$-fold cross-validation
Method Arousal Valence Emotion
ACC (%) std (%) F1 (%) std (%) ACC (%) std (%) F1 (%) std (%) ACC (%) F1 (%)
EEGNet 75.47 17.63 75.54 17.83 73.66 16.82 73.33 16.94 74.57 74.44
DeepConvNet 79.35 15.03 78.43 18.10 72.38 18.87 73.03 19.27 75.87 75.73
ShallowConvNet 80.96 14.49 80.87 14.71 79.77 16.44 80.22 15.60 80.37 80.55
TSception 79.23 11.51 81.30 9.84 78.34 15.03 81.18 12.08 78.79 81.24
EEG-TCNet 76.36 17.05 73.79 19.66 69.72 16.06 54.28 22.43 73.04 64.04
TCNet-Fusion 76.72 17.32 76.95 17.49 75.88 16.70 75.42 16.97 76.30 76,20
ATCNet 82.01 14.06 79.17 17.57 83.19 12.09 82.08 13.41 82.60 80.63
DGCNN 82.12 12.05 81.83 12.45 82.72 11.00 82.40 11.44 82.42 82.12
LGGNet-Fro 81.85 12.57 81.58 12.73 82.34 12.84 81.66 13.91 82.11 81.62
LGGNet-Gen 82.15 12.72 82.15 12.57 82.53 13.20 82.47 13.22 82.34 82.31
LGGNet-Hem 81.92 12.08 81.58 12.43 82.37 12.71 81.58 14.13 82.15 81.58
AT-DGNN-Fro 83.51 11.09 83.06 11.83 85.56 10.69 85.61 10.89 84.54 84.34
AT-DGNN-Gen 83.74 12.03 84.56 10.88 86.01 9.40 85.68 10.06 84.88 85.12
AT-DGNN-Hem 83.73 11.20 84.78 10.93 84.89 10.57 85.35 10.24 84.31 85.07
  • Emotion averages the ACC and F1 scores of both the arousal and valence dimensions.
  • The results for all models are derived from the optimal values among ten trials with distinct random seeds.